OpenStudy (anonymous):

22. A spring is stretched 0.418 m from equilibrium by a force of magnitude 1.00  102 N. (a) What is the force constant of the spring? (b) What is the magnitude of force required to stretch the spring 0.150 m from equilibrium? (c) How much work must be done on the spring to stretch it 0.150 m from equilibrium and to compress it 0.300 m from equilibrium?

OpenStudy (anonymous):

\[F = ks\] where k is the spring constant and s is the distance stretched. (+ positive Tension, - negative compression) so: \[102N = 0.418k\] means that \[k = 102/0.418 = 244 \] b) \[F = 244 * 0.150 = 36.6\] c) \[W = Fd = ks^2\] so \[W_1 = 244*.150^2 = 5.5 J\] and \[W_2 = 244 * (-0.300)^2 = 21.96 = 22 J\]